After graduating with a BCom degree, Tim had interviews with two companies A and B

Suppose that there is a 20% chance that he will be offered the job by company A and a 15%

chance that he will be offered a job with company B. Suppose that each company is unaware

that he has had an interview with the other company. What is the probability that he will be

offered a job by exactly one of the companies? (answer to 2 decimal places):


1
Expert's answer
2022-05-12T18:27:25-0400
"P(exactly\ one)=P(A\cap B^C)+P(A^C\cap B)"

"=0.2(1-0.15)+(1-0.2)(0.15)=0.29"